# Bayesian Unsupervised Disentanglement of Anatomy and Geometry for Deep Groupwise Image Registration

### v2

@@ 5.5.3 Robustness on Different Initial Misalignment
5.5.4 Visualization of Registration Results
@@ 5.7.1 Experimental Design
  • ModeT [106]. We replaced the registration modules of our model by the registration networks proposed in the ModeT paper. Particularly, as both our and the ModeT models have five levels, we follow the same settings in that paper, i.e., for each of the three coarest levels, we first use a ModeT module (Motion Decomposition Transformer) to infer multiple transformations, and then use a CWM (Competitive Weighting Module) to generate a single transformation; for each of the other two levels, we use a ModeT module to infer a single transformation directly.
@@ 5.7.2 Results
Ours-CN Attention U-Net Convs 0.842$\pm$0.051 2.17$\pm$0.58
Ours-CD Attention U-Net Convs 0.836$\pm$0.043 2.21$\pm$0.47
@@
Ours-CD TransMorph Convs 0.848$\pm$0.048 2.09$\pm$0.52
Ours-CN Attention U-Net PIViT 0.841$\pm$0.043 2.18$\pm$0.47
Ours-CD Attention U-Net PIViT 0.838$\pm$0.045 2.26$\pm$0.49
Ours-CN Attention U-Net ModeT 0.848$\pm$0.040 2.12$\pm$0.42
@@ 5.8.1 Structural Representation Maps
These structural representations also demonstrate the clear advantage of our proposed models over conventional learning-based counterparts in terms of model interpretability and registration accuracy. This suggests that the trade-off between model interpretability and performance is not inevitable, while one may believe that more complex models are more accurate. In fact, since medical images are scarce and highly correlated [6], the observations are well structured and only occupy a low-dimensional landscape (or manifold) within the entire data space [107]. Therefore, by respecting the underlying generating process and the inherent symmetry structures, the proposed models achieved better interpretability and higher accuracy, while using fewer parameters.
5.8.2 Latent Symmetry Structures
We also validated the symmetry structures our learnt models should respect through counterfactuals [108, 109]. Fig. 10(a) demonstrates the counterfactual reconstructions by ontological transformations on latent structures. It can be observed that the direct effect of counterfactuals corresponds exactly to the latent structures. Moreover, Fig. 10(b) displays the counterfactual reconstruction by spatial diffeomorphisms on the latent structures. One can see from the difference maps of the equivariance equation that the learnt model is transformation-equivariant. Thus, the proposed model indeed respects the two independent symmetry structures in the observations, and the learnt latent variables are indeed disentangled representations for groupwise registration formally defined in Supplementary Material Section E.
@@ 5.8.2 Latent Symmetry Structures
[…]
@@ 6 Conclusion and Discussion
Furthermore, recent years have witnessed a growing academic interest in the field of interpretable or explainable artificial intelligence (XAI) [121, 107, 122, 123]. Our proposed model exhibits the local and active interpretability characterised by [123] in that the structural representations encoded by the network are deliberately devised to have visual semantics in a human-understandable fashion. The proposed registration module and the network architecture also demonstrate certain levels of algorithmic transparency [107]. Indeed, the Demons algorithm for calculating the velocity fields is mathematically interpretable, as it depends linearly on the gradient and difference of structural representations. On the other hand, the decoder architecture is spatially decomposable, as the reconstruction of image intensities is determined locally by the anatomy at each independent location (owing to the use of convolutions with kernel size 1 isotropically), thus respecting the spatial equivariance in the imaging process w.r.t. diffeomorphic transformations. One of the potential trade-offs or limitations of the proposed architecture is that the usage of convolutions with kernel size 1 in the decoders may provoke a tradeoff between the reconstruction fidelity and the equivariance constraint. Convolutions with kernel size 1 can satisfy the spatial equivariance constraint on the learnt imaging process. However, the lack of network parameters may also lead to suboptimal reconstruction, which can hinder the learning of the registration parameters. Future studies could be addressed to represent the generative process of multi-modal images from the underlying common anatomy in a more compact and unified manner using techniques from, for example, normalising flows [124] that construct an invertible mapping between the observation and the latent variable.
Acknowledgments
Data were provided in part by OASIS-1: Cross-Sectional: Principal Investigators: D. Marcus, R, Buckner, J, Csernansky J. Morris; P50 AG05681, P01 AG03991, P01 AG026276, R01 AG021910, P20 MH071616, U24 RR021382.
@@ Appendix A Minimization of The Intrinsic Distance
[…]
@@ Appendix B The Gumbel-Rao Estimator
where $\bm{G}^{s}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\bm{g}+\log\bm{\pi}\mid\bm{z}$ for $s=1,\dots,S$. Thus, we have the following proposition: let ¯$\overline{\nabla}_{\bm{\psi}}\triangleq\nicefrac{{\operatorname{d}\mathbb{E}[f(\bm{z})]}}{{\operatorname{d}\bm{z}}}$ be the true gradient we are trying to estimate, then
@@ Appendix C The Diffeomorphic Demons Algorithm
$\operatorname*{arg\,max}_{\bm{u}}\frac{1}{2\absolutevalue{\Omega_{F}}}\left\{\norm{\bm{F}-\bm{M}\circ\phi_{s}\circ(\operatorname{id}+\bm{u})}_{F}^{2}+\sigma_{\bm{\varphi}}^{2}\norm{\bm{u}}_{F}^{2}\right\}$ (37)
@@
$E_{s}(\bm{u})$ $=\frac{1}{\absolutevalue{\Omega_{F}}}\sum_{\bm{\omega}\in\Omega_{F}}E_{s}^{\bm{\omega}}(\bm{u})$ (39)
$\triangleq\frac{1}{2\absolutevalue{\Omega_{F}}}\sum_{\bm{\omega}\in\Omega_{F}}\left\{\norm{\bm{\varphi^{\omega}}_{s}(\bm{u})}_{2}^{2}+\sigma_{\bm{\varphi}}^{2}\norm{\bm{u}(\bm{\omega})}_{2}^{2}\right\}$
$\approx\frac{1}{2\absolutevalue{\Omega_{F}}}\sum_{\bm{\omega}\in\Omega_{F}}\norm{\begin{bmatrix}\bm{\varphi^{\omega}}_{s}(\bm{0})\\ \bm{0}\end{bmatrix}+\begin{bmatrix}\bm{J}_{\bm{\varphi}}(\bm{\omega})\\ \sigma_{\bm{\varphi}}\bm{I}_{d}\end{bmatrix}\cdot\bm{u}(\bm{\omega})}_{2}^{2}.$
@@ E.1 Disentangled Representations & Symmetry Groups
    There is an action $:\mathcal{G}\times\mathcal{Z}\rightarrow\mathcal{Z}$;
@@ E.2 Identifiability of Disentangled Representations
If the data generating process follows $u_{t+1}=g\circ m\circ g^{-1}(u_{t})$, then the encoders $\widetilde{g}^{-1}$ that solve the observational identity $\widetilde{g}\circ m\circ\widetilde{g}^{-1}=g\circ m\circ g^{-1}$ identify the true encoder $g^{-1}$ up to equivariances of $m$, i.e. $\widetilde{g}^{-1}\sim_{\mathcal{E}}g^{-1}$ where =$\mathcal{E}=\{a\mid a\text{ is bijective},a\circ m=m\circ a\}$.
@@ F.1 Multi-Modal & Intersubject Groupwise Registration
Figs. 14 to 17 visualise the registration results on a test image group from each dataset using different registration methods. The predicted backward transformation fields are also presented as colourized grid maps, whose Jacobian determinant maps are also visualised. For the proportions of displacement voxels with negative Jacobian, LABEL:tab:ratio_neg_jac_results presents the results for all the four datasets, in addition to the values on the OASIS dataset summarised in the manuscript.
In addition to the visualization of multi-level deformations from our model Ours-CD on the MSCMR dataset, we visualize results from the Learn2Reg, BraTS-2021 and OASIS datasets in LABEL:fig:multilevel_disps_l2r, LABEL:fig:multilevel_disps_brats and LABEL:fig:multilevel_disps_oasis, respectively. We observe a phenomenon similar to MSCMR: the displacements at different levels exhibit different degrees of smoothness, with larger-level $\{\phi_{m}^{l}\}_{m}$ placing greater emphasis on local fine-grained distortion. The deformations $\{\phi_{m}^{1},\phi_{m}^{2}\}_{m}$ for the OASIS image group and the deformation $\{\phi_{m}^{1}\}_{m}$ for the BraTS-2021 group tend to be close to zero, probably because misalignments in these groups primarily occur among local tissues rather than affecting the entire brain structure, as indicated by the original images. In conclusion, these figures of multi-level deformations showcase the effectiveness of our hierarchical decomposition strategy under various imaging conditions.
@@ F.2 Scalability Test on Large-scale and Variable-size Image Groups
LABEL:fig:scalability_app presents the evaluation metrics versus $N_{\text{test}}$ on different datasets.
LABEL:fig:mscmr_scalability demonstrates the result of groupwise registration on a very large image group (120 images from the MS-CMRSeg dataset in total). LABEL:fig:l2r_scalability, LABEL:fig:brats_scalability, LABEL:fig:oasis_scalability further present examples of large-scale groupwise registration results for the Learn2Reg (group size of 16), BraTS (group size of 16) and OASIS (group size of 20) datasets, respectively. One can observe that the proposed model Ours-CD could accurately register all these images, while conventional iterative or learning-based methods can hardly deal with such large group sizes, highlighting the applicability and efficiency of our methods.
LABEL:fig:oasis_age_group displays the mean image of each registered age group in the OASIS dataset using different groupwise registration methods. One can observe that the initial mean images of the unregistered images are fuzzy around the cortical surface. However, after registration using the iterative methods (APE/CTE/$\mathcal{X}$-CoReg), the mean images become sharper, and the proposed models (Ours-CN/Ours-CD) could produce even sharper mean images, especially for relatively large image groups in the regions marked by the red circles.
@@ F.3 Model Interpretability
Fig. 10 presents an example of the learnt structural representation maps for the BraTS data extracted from the models Ours-CN and Ours-CD, respectively. Notably, one can find that complementary regions of the brain are reflected on these representation maps, which are similar across the modalities. Besides, note that these structural representations corresponded spatially to the original images, and thereby are independent of the registration variables. Likewise, Section 5.5.2 presents an example of the learnt structural representation maps for the MS-CMRSeg data from the models Ours-CN and Ours-CD, respectively. One can observe that the representations extracted from Ours-CN may be more semantic and high-level, while those from Ours-CD could be more detailed and local.
@@ F.4 Multi-Modal Counterfactual Reconstructions
F.5 Comparisons of Time Costs

### v3

@@ 5.5.3 Robustness on Different Initial Misalignment
5.5.4 Visualisation of Registration Results
@@ 5.7.1 Experimental Design
  • ModeT [106]. We replaced the registration modules of our model by the registration networks proposed in the ModeT paper. Particularly, as both our and the ModeT models have five levels, we follow the same settings in that paper, i.e., for each of the three coarsest levels, we first use a ModeT module (Motion Decomposition Transformer) to infer multiple transformations, and then use a CWM (Competitive Weighting Module) to generate a single transformation; for each of the other two levels, we use a ModeT module to infer a single transformation directly.
@@ 5.7.2 Results
[…]
@@ 5.8.1 Structural Representation Maps
These structural representations also demonstrate the clear advantage of our proposed models over conventional learning-based counterparts in terms of model interpretability and registration accuracy. This suggests that the trade-off between model interpretability and performance is not inevitable, while one may believe that more complex models are more accurate. In fact, since medical images are scarce and highly correlated [6], the observations are well structured and only occupy a low-dimensional landscape (or manifold) within the entire data space [107]. Therefore, by respecting the underlying generating process and the inherit symmetry structures, the proposed models achieved better interpretability and higher accuracy, while using fewer parameters.
5.8.2 Counterfactual Validation of Latent Symmetries
A deeper level of interpretability can be assessed by examining if the model respects the underlying symmetries of the data-generating process, namely the disentanglement of anatomy and geometry. We validate this by performing counterfactual reconstructions [108, 109], where we actively intervene on the learnt latent variables and observe the outcome. This allows us to ask “what-if” questions and verify that the model behaves in a predictable and semantically meaningful way.
We first investigate the disentangled anatomy representation $\bm{z}$. For a given test image group, we compute the latent anatomy $\bm{z}$ and then perform an ontological intervention by setting a specific channel $\bm{z}_{k}$ to zero (i.e., a $\operatorname{do}(\bm{z}_{k}=\bm{0})$ operation). This corresponds to asking the model: “What would the registered images look like if this specific anatomical feature $k$ did not exist”. The results of this experiment on the OASIS dataset are shown in Fig. 11(a). The figure demonstrates that zeroing out a single channel cleanly removes a specific anatomical structure (e.g., the white matter) from the reconstruction. The difference map between the original and counterfactual reconstructions–the direct effect–clearly isolates this structure. To quantify this observation, we measure the Normalised Cross-Correlation (NCC) between the map of the latent channel $\bm{z}_{k}$ that was set to zero and the corresponding direct effect map. For the example in Fig. 11(a), the four non-zero channels $\bm{z}_{1}$, $\bm{z}_{2}$. $\bm{z}_{4}$ and $\bm{z}_{6}$ yielded high NCC values of 0.99, 0.98, 0.97 and 0.62, respectively. This high correlation confirms that the model has learnt a highly disentangled representation, where individual latent channels have a direct and exclusive correspondence to specific anatomical structures.
This level of granular control and semantic correspondence is a key advantage over conventional end-to-end registration networks. In such “black-box” models, the latent features are typically entangled, and manipulating a single feature is unlikely to produce a coherent or predictable semantic change. Our framework, by explicitly modelling anatomy and geometry as disentangled latent variables, provides interpretable latent that can be used to understand and validate the model’s internal logic, a crucial step towards trustworthy medical AI. We also validate the geometric equivariance in Fig. 11(b), showing that the model correctly disentangles spatial transformations.
@@ 5.8.2 Counterfactual Validation of Latent Symmetries
Figure 11: Counterfactual reconstruction on an image from the OASIS dataset using the underlying symmetry transformations. (a) For ontological transformations acting in the anatomy domain, one can see that the difference calculated by $f(\bm{z};\mathit{do}(\bm{z}_{k}=\bm{0}))-f(\bm{z})$ indeed corresponds to the $k$-th latent structure $\bm{z}_{k}$, which means that the learnt decoder respects the ontological symmetry. (b) For diffeomorphic transformations acting in the spatial domain, one can observe that the equivariance difference $f(\bm{z}\circ\phi_{i})-f(\bm{z})\circ\phi_{i}$, where $f(\bm{z}\circ\phi_{i})=f(\bm{z}\circ\phi;\mathit{do}(\phi=\phi_{i}))$, are almost zero except for interpolation errors, which indicates that the learnt decoder is indeed transformation-equivariant.
@@ 6 Conclusion and Discussion
Furthermore, recent years have witnessed a growing academic interest in the field of interpretable or explainable artificial intelligence (XAI) [121, 107, 122, 123]. Our proposed model exhibits the local and active interpretability characterised by [123] in that the structural representations encoded by the network are deliberately devised to have visual semantics in a human-understandable fashion. The proposed registration module and the network architecture also demonstrate certain levels of algorithmic transparency [107]. Indeed, the Demons algorithm for calculating the velocity fields is mathematically interpretable, as it depends linearly on the gradient and difference of structural representations. On the other hand, the decoder architecture is spatially decomposable, as the reconstruction of image intensities is determined locally by the anatomy at each independent location (owing to the use of convolutions with kernel size 1 isotropically), thus respecting the spatial equivariance in the imaging process w.r.t. diffeomorphic transformations. One of the potential trade-offs or limitations of the proposed architecture is that the usage of convolutions with kernel size 1 in the decoders may provoke a trade-off between the reconstruction fidelity and the equivariance constraint. However, by sacrificing a small degree of reconstruction fidelity, we gain a significant improvement in the identifiability and accuracy of the geometric transformations. This is an example of how imposing a strong, domain-appropriate inductive bias can guide a learning system to a more robust and accurate solution for its primary task. Future studies could be addressed to represent the generative process of multi-modal images from the underlying common anatomy in a more compact and unified manner using techniques from, for example, normalising flows [124] that construct an invertible mapping between the observation and the latent variable.
@@ Appendix A Minimization of The Intrinsic Distance
$\left\{\begin{aligned} \frac{1}{N}\sum_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}&=\frac{\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}{\sum_{k=1}^{K}\left[\prod_{j=1}^{N}\bm{\pi}_{j,k}^{\diamond}\right]^{\nicefrac{{1}}{{N}}}}&&k=1,\dots,K\\ \sum_{k=1}^{K}\bm{\pi}_{j,k}^{\diamond}&=\bm{1},&&j=1,\dots,N\end{aligned}\right$. (27)
@@ Appendix B The Gumbel-Rao Estimator
where $\bm{G}^{s}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}\bm{g}+\log\bm{\pi}\mid\bm{z}$ for $s=1,\dots,S$. Thus, we have the following proposition: let $\overline{\nabla}_{\bm{\psi}}\triangleq\nicefrac{{\operatorname{d}\mathbb{E}[f(\bm{z})]}}{{\operatorname{d}\bm{z}}}$ be the true gradient we are trying to estimate, then
@@ Appendix C The Diffeomorphic Demons Algorithm
[…]
@@
[…]
@@ E.1 Disentangled Representations & Symmetry Groups
[…]
@@ E.2 Identifiability of Disentangled Representations
If the data generating process follows $u_{t+1}=g\circ m\circ g^{-1}(u_{t})$, then the encoders $\widetilde{g}^{-1}$ that solve the observational identity $\widetilde{g}\circ m\circ\widetilde{g}^{-1}=g\circ m\circ g^{-1}$ identify the true encoder $g^{-1}$ up to equivariances of $m$, i.e. $\widetilde{g}^{-1}\sim_{\mathcal{E}}g^{-1}$ where $\mathcal{E}=\{a\mid a\text{ is bijective},a\circ m=m\circ a\}$.
@@ F.1 Multi-Modal & Intersubject Groupwise Registration
Figs. 14 to 17 visualise the registration results on a test image group from each dataset using different registration methods. The predicted backward transformation fields are also presented as colorized grid maps, whose Jacobian determinant maps are also visualized. For the proportions of displacement voxels with negative Jacobian, Table V presents the results for all the four datasets, in addition to the values on the OASIS dataset summarised in the manuscript.
In addition to the visualization of multi-level deformations from our model Ours-CD on the MSCMR dataset, we visualize results from the Learn2Reg, BraTS-2021 and OASIS datasets in Fig. 18, Fig. 19 and Fig. 20, respectively. We observe a phenomenon similar to MSCMR: the displacements at different levels exhibit different degrees of smoothness, with larger-level $\{\phi_{m}^{l}\}_{m}$ placing greater emphasis on local fine-grained distortion. The deformations $\{\phi_{m}^{1},\phi_{m}^{2}\}_{m}$ for the OASIS image group and the deformation $\{\phi_{m}^{1}\}_{m}$ for the BraTS-2021 group tend to be close to zero, probably because misalignments in these groups primarily occur among local tissues rather than affecting the entire brain structure, as indicated by the original images. In conclusion, these figures of multi-level deformations showcase the effectiveness of our hierarchical decomposition strategy under various imaging conditions.
@@ F.2 Scalability Test on Large-scale and Variable-size Image Groups
Fig. 21 presents the evaluation metrics versus $N_{\text{test}}$ on different datasets.
Fig. 22 demonstrates the result of groupwise registration on a very large image group (120 images from the MS-CMRSeg dataset in total). Fig. 23, Fig. 24, Fig. 25 further present examples of large-scale groupwise registration results for the Learn2Reg (group size of 16), BraTS (group size of 16) and OASIS (group size of 20) datasets, respectively. One can observe that the proposed model Ours-CD could accurately register all these images, while conventional iterative or learning-based methods can hardly deal with such large group sizes, highlighting the applicability and efficiency of our methods.
[…]
@@ F.3 Model Interpretability
Fig. 27 presents an example of the learnt structural representation maps for the BraTS data extracted from the models Ours-CN and Ours-CD, respectively. Notably, one can find that complementary regions of the brain are reflected on these representation maps, which are similar across the modalities. Besides, note that these structural representations corresponded spatially to the original images, and thereby are independent of the registration variables. Likewise, Fig. 28 presents an example of the learnt structural representation maps for the MS-CMRSeg data from the models Ours-CN and Ours-CD, respectively. One can observe that the representations extracted from Ours-CN may be more semantic and high-level, while those from Ours-CD could be more detailed and local.
@@ F.4 Multi-Modal Counterfactual Reconstructions
F.5 Comparisons of Time Costs of Baseline and Proposed Methods
@@ F.5.3 Time Costs for Achieving 95% of the Best Performance
F.6 Comparisons of Time Costs of Proposed Method Integrated With State-of-the-Art Registration Backbones
F.6.1 Training Time
In Table VII we report the training time required to reach (i) the best Dice score (DSC) of each method, (ii) 95% of the best DSC, and (iii) the best DSC achieved by the strongest baseline method (CTE-Att).
The results indicate that our model maintains competitive training efficiency when integrated with state-of-the-art registration backbones. In particular, integration with TransMorph leads to significantly shorter training time for reaching either 95% of the best DSC or the best baseline DSC, highlighting its superior convergence speed. The total training time is also comparable to our original Att-UNet version, while achieving better registration performance. Similar trends are observed when using ModeT as the backbone.
The only substantial increase in training time occurs with PIViT. We would like to note that this overhead is not inherent to our model. During experiments with PIViT, we observed training instability, which was mitigated by reducing the learning rate from $10^{-3}$ to $10^{-4}$. While this change stabilized convergence, it also led to a longer training process. Notably, such instability was not encountered with TransMorph or ModeT, suggesting that the issue is more likely attributable to PIViT’s internal training dynamics rather than any limitation in our integration mechanism.
F.6.2 Inference Time
In Table VIII, we compare the inference time of the original model and its integrated versions. It can be observed that our model achieves very fast inference speed, requiring only about 0.04 seconds per image group from the MS-CMRSeg dataset. Remarkably, integrating our model with state-of-the-art registration backbones does not introduce any inference time overhead. This demonstrates the efficiency and practicality of our framework for real-time medical image processing. Furthermore, the consistent inference speed across different integration settings highlights the modularity of our method, making it readily adaptable to various registration pipelines without compromising performance.
Overall, these results demonstrate that our framework can be integrated with various state-of-the-art registration backbones without introducing substantial training or inference time overhead. This confirms that our method is both flexible and computationally efficient in practice.
F.7 Ablation Study of Reconstruction Fidelity
We performed ablation experiments using Ours-CD on the MS-CMRSeg dataset to quantify the trade-off of reconstruction fidelity and registration performance. The results are shown in Table IX.
First, we replaced the decoder kernel size from 1 to 3. The results showed that kernel size 1 achieved slightly higher average DSC and ASSD compared to kernel size 3, confirming that the minor loss in reconstruction fidelity does not compromise registration performance. Considering kernel size 3 may introduce theoretical issues as discussed earlier, choosing kernel size 1 is thus well justified to ensure theoretical soundness without compromising performance.
Second, we conducted experiments where we kept the decoder kernel size at 1, but progressively reduced the weight of the reconstruction loss. We observed that as the reconstruction loss weight decreased, the registration performance also declined. This confirms that while photorealistic reconstruction is not the goal, the reconstruction loss provides a critical learning signal that supports accurate registration.
In conclusion, the trade-off is real, but it is one we make deliberately and advantageously. By sacrificing a small degree of reconstruction fidelity, we gain a significant improvement in the identifiability and accuracy of the geometric transformations. This is an example of how imposing a strong, domain-appropriate inductive bias can guide a learning system to a more robust and accurate solution for its primary task.
F.8 Ablation Study of Hierarchical Level $L$
We conducted a parameter study for $L$ while keeping all other hyperparameters fixed. Specifically, we trained Ours-CD with $L=3,4,5,6$ on the MS-CMRSeg dataset, and report the corresponding average Dice scores (DSC) on the test set:
  • $L=3$: 0.782
  • $L=4$: 0.802
  • $L=5$: 0.836
  • $L=6$: 0.801
These results demonstrate that our default choice of $L=5$ yields the best performance, with both smaller and larger $L$ values leading to a noticeable drop.
This trend can be attributed to several interrelated factors:
  1. 1.
    Architectural Alignment with Hierarchical Inference. In our current implementation, the number of levels $L$ is tightly coupled with the depth of the network (based on an Attention U-Net). A smaller $L$ results in a shallower network with fewer downsampling stages, reducing its representational capacity and limiting its ability to model complex deformations—leading to underfitting. Conversely, increasing $L$ adds deeper layers and coarser levels, which do not always translate to better performance, because larger $L$ values introduce additional layers and parameters, increasing the risk of overfitting and making optimization more challenging.
  2. 2.
    Overly Coarse Deformations at Large $L$. At $L=6$, the coarsest feature map operates at a spatial resolution of only $5\times 5$ (for a $160\times 160$ input). The velocity field at this scale is tasked with modeling extremely global deformations, which are not always necessary and may introduce instability or optimization noise. In practice, such coarse scales often fail to yield meaningful gradient signals.
  3. 3.
    Diminishing Returns Beyond Optimal Hierarchical Depth. The hierarchical decomposition is most effective when each scale contributes interpretable sub-fields corresponding to different spatial frequencies. When $L$ becomes too large, the added coarse levels may no longer correspond to meaningful anatomical structures, reducing the efficacy of the decomposition and potentially disrupting optimization.
We would also like to emphasize that these limitations are not inherent to the proposed framework itself, but rather a by-product of our current architectural implementation. In principle, the model design allows for more flexible architectures that decouple $L$ from network depth. For example:
  • One could avoid strict $2\times$ downsampling at each level, maintaining higher spatial resolutions at deeper levels;
  • One could share network parameters across levels to avoid increasing the parameter count with $L$.
We believe that these modifications could allow for a further scaling of $L$ without the drawbacks observed here, and we plan to explore architectural variants that better adapt to varying $L$ in future work.